Q3) The following data is given;

Number of Employed = 33000

Labor Force = 36000

Working Age Population = 54000

  1. Calculate the unemployment rate
  2. Calculate the labor force participation rate
  3. Calculate the employment to population ratio.
Answer #1

a) Unemployment rate = Total unemployed * 100 / total labour force

Total unemployed = Labour force - number of employed = 36000 - 33000 = 3000

Unemployment rate = 3000 * 100 /36000 = 8.33%.

b) Labour participation rate = Labour force * 100 / working age population = 36000 * 100 / 54000 = 66.66%

c) Employment to population ratio = Number of employed / total population

Here total population is not given so we could use working age population as total popuation.

Employment to population ratio = 33000/54000 = 0.611

#please rate positively/\...thank you.